OpenStudy (anonymous):

(2x-5)(4x^2-3x+1) giving out medal

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

i think you want to simplify it. \[\left( 2x-5 \right)\left( 4x^2-3x+1 \right)=2x(4x^2-3x+1)-5(4x^2-3x+1)\] \[=8x^3-6x^2+2x-20x^2+15x-5=?\]